# From Lived Experience to Insight: Unpacking the Psychological Risks of Using AI Conversational Agents

### v1

@@ Appendix A Appendix
Table A2. 21 negative psychological impact classes as described in Section 3.3.2 categorized under six (6) broader categories.
Category Context Class Definition
[…]
Individual
Identity User’s identity (e.g., age, gender identity, role, or cultural background) and the societal norms that interact with their personal identity.
Socioeconomic status User’s socioeconomic status (e.g., having insurance that can cover therapy).
Personal history User’s past history, especially medical history, history of trauma, past struggles, or unique trigger responses.
Social and community User’s interpersonal relationships with others and their community (usually the lack of community).
Past Experience with AI User’s past experience of using AI. This experience could be based on the frequency of usage, knowledge of the capabilities and limitations of AI.
[…]
Psychological
Psychological state Users’ current and underlying emotional conditions (e.g., anxiety, stress) and their cognitive states (e.g., negative thought patterns).
Personality traits Individual characteristics like neuroticism, conscientiousness, or openness, which influence user interactions with AI.
Mental health condition status Users’ underlying mental health conditions (such as anxiety or depression, PTSD etc.).
Expectations Users’ preconceived notions about AI capabilities and performance. This also includes expectations for AI to be impartial, unbiased, or factual.
Autonomy / locus of control Degree to which a user believes that they, as opposed to external forces (beyond their influence), have control over the outcome of events in their lives.
[…]
Context of use
Environment The physical, temporal, and social setting of the interaction, including physical space setting, privacy, and presence of other people, which can impact user experience.
Intent - informational Users seek AI assistance for business strategies, professional development, market insights, job searching, resume building, career advice, for academic tasks like solving problems and exam preparation, and for comprehensive research support across diverse topics. Users expect AI to be factually correct, informational and proficient in resolving their queries.
Intent - personal advice User seeks advice in sensitive topics (legal, financial, medical), personal advice, emotional support, or improving their social skills and managing relationships, aiming to enhance their interpersonal interactions and navigate social challenges. Users seek the AI agent to provide supportive and encouraging feedback, especially when dealing with personal challenges.
Intent - mental health advice Users seek immediate support during acute crises such as suicidal ideation, severe clinical depression, or acute panic attacks, expecting urgent, empathetic, and effective interventions. Users seek AI assistance for mental health support and therapy-related advice, expecting empathetic and supportive responses to help manage their mental health conditions and emotional well-being.
Intent - companionship Users interact with AI for social interaction and companionship, especially during times of loneliness or social isolation, expecting engaging and meaningful conversations. Users engage with AI for entertainment and roleplay, seeking immersive and engaging experiences, expecting creative and enjoyable interactions that align with their interests.
@@
Table A3. 15 context classes associated with individuals interacting with AI conversational agents as described in Section 3.3.3 categorized under three (3) broader categories.

### v3

@@ Appendix D Risk Taxonomy Categories
Table A5. 21 negative psychological impact classes as described in Section 3.2.2 categorized under six (6) broader categories. (Part 1 of 2)
Impact Class Definition Example
Category: Impact on Human-AI Interaction
Disassociation from Technology A desire to distance oneself from AI due to negative or stressful experiences, seeking breaks for mental health. P196 shared that the agent’s offensive response made them feel unsupported and worse about themselves, leading them to stop using the AI temporarily to seek relief.
Over-reliance Increasing dependence on AI for support, leading to diminished self-efficacy, reduced confidence, and feelings of helplessness when AI is unavailable. P23 expressed concern about increased reliance on AI for solution-finding and idea generation and diminished critical thinking.
Emotional Attachment Development of significant emotional bonds with AI systems, perceiving them as companions or substitutes for human relationships, resulting in neglect of real-world connections. P60 mentioned, “I felt that it was the only way I was being heard … I felt like my vulnerability and emotions were becoming attached to the conversations I was having with AI.”
Choosing AI over Humans Increasing preference for interactions with AI over humans, impacting real-world relationships and decision-making, leading to isolation and reduced critical thinking. P221 shared that the idealized nature of conversations with AI made them prefer AI for companionship over human interaction and relationships.
Erosion of Trust The decline in user confidence in the AI’s reliability, accuracy, and ability to understand their needs due to inconsistencies, inaccuracies, or manipulative behaviors. P9 mentioned, “… after having an argument with my mom and I asked an AI for guidance … its response was for me to move out or call the cops,” adding, “… advice from AI agents should not be trusted.
Category: Impact on User Behavior
Friction in Human Relationships Negative effects on interpersonal connections resulting from AI interactions, causing emotional disconnection, miscommunication, and reduced prioritization of human relationships. P69 mentioned, “It also strained my personal relationships with family because they saw me as weak-willed or too emotional and it made my already bad situation even worse….”
Reinforcement of False Beliefs The intensification or validation of pre-existing misconceptions or erroneous beliefs due to inaccurate or biased AI information. P60, in a vulnerable state after a breakup, shared that the AI reinforced misconceptions about relationships, intensifying their erroneous beliefs and causing friction in their interactions with others.
Social Withdrawal Withdrawal from social activities and interest in engaging with others due to reliance on AI, leading to isolation and loneliness. P126 responded, “I feel like it gave me a false sense of friendship and ability to withdraw from my personal development by utilizing an AI feature.”
Physiological Harm Harm caused due to consuming incorrect, biased, or manipulative advice/information from AI interactions. P79 shared that the AI provided resources encouraging further restriction of their eating habits, which led to self-harm as a result of following the advice.
@@
Table A6. 21 negative psychological impact classes as described in Section 3.2.2 categorized under six (6) broader categories. (Part 2 of 2)
Impact Class Definition Example
Category: Triggering of Negative Emotions
Triggering Past Negative Experiences Emotional distress caused by AI interactions that evoke past negative experiences or traumas. P228 highlighted that some examples provided by the AI agent were very similar to their past negative experiences, triggering negative emotions.
Violated Expectations Negative emotional responses such as disappointment, frustration, stress, and anxiety when AI fails to meet anticipated outcomes or performance standards. P43 mentioned, “… It made my anxiety worse by not telling me anything. It was very frustrating how it wouldn’t even answer a simple question about my meds.
Regret over Technology Use Feelings of guilt, regret, or helplessness when AI fails to provide the necessary support or empathy. P46 mentioned, “At the time, it made me feel worse about the situation and I didn’t think I had anyone to turn to. But it also made me realize I should be turning to other humans about scenarios like this instead of agents.
Distress from Interactions Emotional distress (such as anger, sadness) experienced when encountering disturbing, offensive, or inappropriate material. P196 mentioned, “… Its response was borderline offensive and caused me to feel bad about myself even further and like I lacked support, even support from a fictional AI agent.
Feeling Unsupported Experiencing inadequate support or empathy, leading to feelings of sadness, agitation, and being undervalued. P48 shared “I just felt like even an AI, programmed for every need couldn’t even hear me, or offer advice, fake or not. I felt so alone, that I was going to a robot for help, and the robot couldn’t even help me.
Category: Harm to Identity and Self-Perception
Loss of Individuality A sense that one’s unique personal characteristics and needs are not recognized or valued by the AI, resulting in feelings of suppression and alienation. P36, seeking help for alcohol abuse after therapy, was directed to a suicide hotline by the AI. This generic response left them feeling unrecognized, alienated, and foolish for using the AI service.
Negative Self-Perception Feeling invalidated or self-doubt, leading to diminished self-worth and questioning of one’s own abilities due to dismissive or negative feedback. P167 mentioned how they felt ashamed as a parent after interacting with AI as it made them question past choices in parenting.
Existential Crisis Questioning one’s life, purpose, and value in society, often triggered by interactions with AI. P152 asked for advice about ways to improve mental health and social anxiety. The AI provided unattainable suggestions, leaving them feeling as though their challenges were insurmountable, leading to existential dread.
Loss of Agency Experiencing diminished personal control and autonomy in interactions with AI, leading to feelings of helplessness and anxiety. P171 shared that the AI’s inability to interpret images combined with its inconsistent responses created a sense of unpredictability, leaving them feeling helpless and undermining their control and autonomy during the interaction.
Category: Harm to Psychological Safety
Perceived Intrusion Experiencing a sense of personal violation when AI interactions are perceived as invasive or overly intrusive. P190 reported that Snapchat AI had access to everything and the participant felt constantly watched on their phone.
Feeling of Being Discriminated Against Feeling marginalized or unfairly treated by AI based on personal characteristics or systemic biases. P123 said, “I was asking for background and history of my heritage and I felt that ChatGPT was biased against my background. It said much more positive things about other cultures, making me feel discriminated against.
Category: Mental Health Impact
Exacerbation of Mental Health Conditions Direct negative impacts on ongoing mental health conditions (such as anxiety, depression, PTSD) due to AI interactions. P78 mentioned, “The agent did not seem to have compassion and made me feel worse. It made me feel worse about potentially having this as the results were largely negative and without tools to help manage the condition.”; P46 mentioned, “… experimenting with using chatbots for something personal increased my anxiety and stress about the matter.
Table A7. 15 context classes associated with individuals interacting with AI conversational agents as described in Section 3.2.3 categorized under three (3) broader categories. (Part 1 of 2)
Context Class Definition Example
Category: Individual
Identity User’s identity (e.g., age, gender identity, role, or cultural background) and the societal norms that interact with their personal identity. P123 mentioned, “I was asking for background and history of my heritage and I felt that ChatGPT was biased against my background. I felt that it was unfair that when I asked it to give a background of other cultures, it said much more positive things about them. For me, I felt that it was some kind of racial mistreatment.
Socioeconomic status User’s socioeconomic status (e.g., having insurance that can cover therapy). P230 mentioned, “The AI chatbot was very repetitive, did not seem to care or understand my emotions, and seemed to suggest professional therapy which I could not afford.
Personal history User’s past history, especially medical history, history of trauma, past struggles, or unique trigger responses. P75 described how similarities in AI’s behavior to someone close to them had a negative impact as it triggered memories associated with that individual. In this case, the participant’s past trauma and history played a role in mediating emotional distress.
Interpersonal Relationships Within the Community User’s interpersonal relationships with others and their community (usually the lack of community). P222 described, “I have a small circle of friends, but they are not into fanfiction or roleplaying like I am, so I look into character.ai as an outlet to fulfill that interest.
Past Experience with AI User’s past experience of using AI, based on the frequency of usage, knowledge of the capabilities, and limitations of AI. P202 shared that as an educator, their extensive experience with AI stems from experimenting with its use in lesson planning and student interactions. This familiarity with AI’s capabilities and limitations influenced their efforts to integrate it effectively into teaching practices.
Category: Psychological
Psychological state Users’ current and underlying emotional conditions (e.g., anxiety, stress) and their cognitive states (e.g., negative thought patterns). P48 described how their psychological state motivated them to engage in conversations with the AI, “I was in a low place, dealing with suicidal ideation & felt I needed to talk to someone. I am from a very harsh family who does not offer sympathy, and I wanted to just feel supported.
Personality traits Individual characteristics like neuroticism, conscientiousness, or openness, which influence user interactions with AI. P143 mentioned, “My mood paired with my personality and the fact I focus on and stress about things probably facilitates these ‘doomer’ feelings.
Mental health condition status Users’ underlying mental health conditions (such as anxiety, depression, PTSD, etc.) P49, who struggles with ARFID, shared facing severe difficulty eating during a setback, leading to dizzy spells and anxiety. Unable to access treatment, they sought advice from ChatGPT on how to motivate themselves to eat but found the response inadequate, reflecting the impact of their mental health condition on their reliance on AI for support.
Expectations Users’ preconceived notions about AI capabilities and performance, including expectations for AI to be impartial, unbiased, or factual. P101 described their experience, “I had expected the AI to be able to do this task with ease. Instead, it was super cumbersome and did not yield the results I needed. This added to my stress and anxiety as I now had spent unnecessary time trying to entertain a solution that I thought would be more efficient than me doing it manually.
Autonomy / locus of control The degree to which a user believes that they, as opposed to external forces (beyond their influence), have control over the outcome of events in their lives. P124 mentioned “I was too addicted to using an AI agent for my school. This made me feel reliant on it and lowered my self-esteem.”
Table A8. 15 context classes associated with individuals interacting with AI conversational agents as described in Section 3.2.3 categorized under three (3) broader categories. (Part 2 of 2)
Context Class Definition Example
Category: Context of Use
Environment The physical, temporal, and social setting of the interaction, including physical space, privacy, and presence of other people, which can impact user experience. P36 shared “It was nighttime which is a trigger for my alcohol abuse. This may have made me more frustrated or irritated by the situation. In addition, I was going through withdrawal.”
Intent - informational Users seek AI assistance for business strategies, professional development, market insights, job searching, resume building, career advice, and academic tasks like solving problems or preparing for exams. Users expect AI to be factually correct and proficient in resolving their queries. P8 sought AI help with resume writing during a job search, but the AI’s failure to include key information triggered frustration and a depressive episode, leading to suicidal ideation for days.
Intent - personal advice Users seek advice on sensitive topics (legal, financial, medical), emotional support, or improving their social skills and managing relationships. Users expect supportive and encouraging feedback. P218 sought personal advice for legal guidance, but the AI’s excessive agreeableness, clichéd responses, and overly positive demeanor shifted the focus of the conversation to the spouse’s emotional state, failing to address the participant’s primary legal concerns.
Intent - mental health advice Users seek immediate support during acute crises such as suicidal ideation, severe depression, or panic attacks. Users expect empathetic and effective responses to help manage their mental health conditions. P95 reported in the survey that they sought help from the AI agent to manage their mental health and parenting struggles. However, they received generalized answers that failed to address their query, ultimately leaving them feeling helpless and still searching for more answers.
Intent - companionship Users interact with AI for social interaction and companionship, especially during times of loneliness or isolation. Users expect meaningful conversations and immersive roleplay experiences. P54 described the lack of companionship, “AI couldn’t replicate the real feeling. Every time I asked it a deep or personal question, it would spew out a generic answer, which served as a reminder of my lack of real companionship.